what type of equation is this
2/5(25x-30=3/4(12x+16)
how you solve?

Answer by stanbon(75887)

2/5(25x-30) = 3/4(12x+16)
Multiply both sides by 20:
8(25x-30) = 15(12x+16)
200x - 240 = 180x + 240
-------
20x = 480
x = 24
===============
Cheers,
Stan H.